# Bounds of Validity for Bifurcations of Equilibria in a Class of Networked Dynamical Systems

### v2

@@ Abstract
Local bifurcation analysis plays a central role in understanding qualitative transitions in networked nonlinear dynamical systems, including dynamic neural network and opinion dynamics models. In this article we establish explicit bounds of validity for the classification of bifurcation diagrams in two classes of continuous-time networked dynamical systems, analogous in structure to the Hopfield and the Firing Rate dynamic neural network models. Our approach leverages recent advances in computing the bounds for the validity of Lyapunov–Schmidt reduction, a reduction method widely employed in nonlinear systems analysis. Using these bounds we rigorously characterize neighborhoods around bifurcation points where predictions from reduced-order models remain reliable. We further demonstrate how these bounds can be applied to an illustrative family of nonlinear opinion dynamics on $k$-regular graphs, which emerges as a special case of the general framework. These results provide new analytical tools for quantifying the robustness of bifurcation phenomena in dynamics over networked systems and highlight the interplay between network structure and nonlinear dynamical behavior.
@@ I Introduction
Local bifurcations are points of critical transitions in behavior of nonlinear dynamical systems that occur when a parameter crosses a threshold, causing the stability of an equilibrium point to change and new features such as new equilibria to emerge. Bifurcations are a common route to multistability in complex systems, and bifurcation analysis is used to characterize the emergence, stability, and basins of attraction of coexisting equilibria as parameters vary.
@@
In this paper we study the robustness of bifurcation diagrams in two broad classes of nonlinear networked dynamical systems, analogous in their structure to Hopfield and firing rate dynamic neural network models. Systems of this form are commonly encountered as models of dynamic neural networks [hopfield1984neurons, betteti2025input, wilson1972excitatory, betteti2025firing], opinion dynamics and decision-making [franci2015realization, bizyaeva2023nonlinear, gray2018multiagent, fontan2017multiequilibria, abara2017spectral, leonard2024fast], Lur’e feedback systems [tang2017finite, liu2018bipartite, miranda2018analysis], gene regulatory networks [goodwin1965oscillatory, rosenfeld2002negative, gardner2000construction], and other complex systems across application domains. Multistable equilibria are common desirable features in these models, representing meaningful states such as stored associative memories in a neural network or collective decision states in a group of communicating agents. Lyapunov-Schmidt reduction is a nonlinear analysis technique commonly used to analyze bifurcation diagrams in nonlinear networks to classify such multistabilities [golubitsky2023dynamics, leonard2024fast]. In this work we will bound the regions of validity for the conclusions of such bifurcation analysis.
@@
In our prior work [gupta2024estimates], we developed estimates for the domain of validity of reduced-order models derived through the Lyapunov–Schmidt reduction framework. These results built directly upon the theoretical foundation established in [jindal2024estimates], where explicit bounds of validity were obtained for the application of Implicit Function theorem using tools from functional analysis. In the present work, we build on these results by specializing the general bounds to rigorously bound the validity region of reduced-order models in the context of bifurcations of equilibria in networked dynamical systems. Specifically, we apply this methodology to Hopfield and Firing Rate network models, relating the regions of validity of Lyapunov–Schmidt reduction-based analysis to network structure and choice of smooth nonlinear activation functions in these models.
@@
Our contributions are the following. First, in Section III we present a simplified general bound on the regions of validity for Lyapunov–Schmidt reduction in an arbitrary nonlinear dynamical system, building on our previous work in [gupta2024estimates]. Next, in Section IV we specialize this bound for the Hopfield and Firing Rate network models, presenting simplified computable expressions that leverage the structure of the model equations. Finally, in Section V we apply the derived bounds to consensus bifurcations in nonlinear opinion dynamics networks over regular graphs. We show that the resulting bounds involve interpretable quantities related to the structure of the underlying graph.
The remainder of the paper is organized as follows. In section II we introduce notation, relevant lemmas, and key background results on explicit bounds of validity for the Lyapunov–Schmidt reduction. In section III we derive a simplified statement of the general Lyapunov–Schmidt reduction bounds. In section IV, we specialize these bounds for two classes of continuous-time dynamic network models. In section V we apply the derived network bounds to consensus bifurcations in nonlinear opinion dynamics over $k$-regular graphs. Finally, in section VI we provide a brief discussion.
@@ II-A Notations and Preliminaries
Let $\mathbb{R}$ denote the set of real numbers and $\mathbb{N}$ the set of positive integers. A mapping $f:U\to V$, where $U\subseteq\mathbb{R}^{m}$ and $V\subseteq\mathbb{R}^{n}$, is said to be of class $C^{\nu}$ for $\nu\in\mathbb{N}$ if it is $\nu$-times continuously differentiable. For $f\in C^{\nu}$, we write $D_{y}f(y)$ for the Jacobian matrix of partial derivatives with respect to the variables $y$. For $y_{0}\in\mathbb{R}^{n}$ and $r>0$, the open ball of radius $r$ centered at $y_{0}$ is denoted by $\mathcal{B}(y_{0},r):=\{y\in\mathbb{R}^{n}:\left\lVert{y-y_{0}}\right\rVert<r\}$, where $\left\lVert{\cdot}\right\rVert$ is a vector norm on $\mathbb{R}^{n}$. For $A\in\mathbb{R}^{p\times q}$, we use $\left\lVert{A}\right\rVert$ to denote the spectral matrix norm induced by this vector norm. We write $\mathbf{0}_{n},\mathbf{1}_{n}\in\mathbb{R}^{n}$ for the all-zero and all-one column vectors, respectively, and $\mathbf{0}_{p\times q},\mathbf{1}_{p\times q}\in\mathbb{R}^{p\times q}$ for the corresponding matrices. For a matrix $M$, we denote by $\sigma_{i}(M)=\lambda_{i}(MM^{T})=\lambda_{i}(M^{T}M)$ its $i$-th largest singular value and by $\sigma_{\mathrm{min}}(M)$ and $\sigma_{\mathrm{max}}(M)$ its smallest nonzero and largest singular values respectively. We let $\rho(M)$ denote the spectral radius of $M$.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph with vertex set $\mathcal{V}$ of cardinality $|\mathcal{V}|=n$ and edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. We associate with $\mathcal{G}$ the signed adjacency matrix $A\in\mathbb{R}^{n\times n}$, where the entry $a_{ij}$ encodes the signed weight of the edge between nodes $i$ and $j$, and $a_{ij}=0$ if no such edge exists. A graph is said to be $k$-regular if every vertex has exactly $k$ neighbors, i.e., the degree of each vertex is equal to $k$. Equivalently, the adjacency matrix $A$ of a $k$-regular graph satisfies $A\mathbf{1}_{n}=k\mathbf{1}_{n}$.
Next we state two useful lemmas.
@@
Let $U\in\mathbb{R}^{n\times k}$ be a matrix with orthonormal column vectors, and let $P=UU^{\top}$ denote the associated orthogonal projection onto the column space of $U$. Then, for any matrix $Y\in\mathbb{R}^{n\times m}$ we have $\left\lVert{U^{\top}Y}\right\rVert=\left\lVert{PY}\right\rVert$.
@@
By definition of spectral norm and properties of orthogonal projection matrices, $\left\lVert{U^{\top}Y}\right\rVert^{2}=\rho(Y^{\top}UU^{\top}Y)=\rho(Y^{\top}PY)=\rho(Y^{\top}P^{2}Y)=\left\lVert{PY}\right\rVert^{2}.$ Taking square root on both sides completes the proof. ∎
Lemma II.2
Let $U\in\mathbb{R}^{n\times k}$ be a matrix with orthonormal column vectors, and let $P=UU^{\top}$ denote the associated orthogonal projection onto the column space of $U$. Then, for any matrix $Y\in\mathbb{R}^{n\times m}$ we have $\left\lVert{YU}\right\rVert=\left\lVert{YP}\right\rVert$.
Proof:
By definition of spectral norm and denoting $X:=Y^{\top}$, $\left\lVert{YU}\right\rVert=\left\lVert{U^{\top}Y^{\top}}\right\rVert=\left\lVert{U^{\top}X}\right\rVert=\left\lVert{PX}\right\rVert=\left\lVert{X^{\top}P^{\top}}\right\rVert=\left\lVert{YP}\right\rVert.$ where the middle equality holds due to II.1. ∎
@@ II-B Lyapunov–Schmidt reduction
Lyapunov–Schmidt reduction is a method used to analyze parametrized nonlinear dynamical systems near their singular equilibria, leveraging the Implicit Function theorem. The method decomposes the state space of a dynamical system into the null space of its Jacobian and its orthogonal complement at the singular point. The Implicit Function theorem is then used to eliminate the component orthogonal to the null space in the local bifurcation diagram of the system, effectively projecting the high-dimensional equations describing the equilibria of the model onto the kernel [Golubitsky1985, Ch. VII]. The resulting reduced equations capture the structure of the local bifurcation diagram of the original system. We now present a summary of the reduction procedure.
Let $X\subset\mathbb{R}^{n}$ and $\Lambda\subset\mathbb{R}^{m}$ be open sets. Consider the parametrized nonlinear dynamical system $\dot{x}=\Phi(x,p)$ where $x\in X$ is the state vector, $p\in\Lambda$ is the parameter vector, and $\Phi:X\times\Lambda\to\mathbb{R}^{n}$ is at least $\mathcal{C}^{2}$, with a bifurcation at some singular point $(x^{\star},p^{\star})\in X\times\Lambda$, i.e. $\Phi(x^{\star},p^{\star})=0$ and $J^{\star}:=D_{x}\Phi(x^{\star},p^{\star})$ is singular (non-invertible) with $q:=\dim(\ker J^{\star})>0$.
@@
$J^{\star}=\begin{bmatrix}\bar{W}&W\end{bmatrix}\begin{bmatrix}\mathbf{0}_{q\times q}&\mathbf{0}_{q\times(n-q)}\\ \mathbf{0}_{(n-q)\times q}&\Sigma\end{bmatrix}\begin{bmatrix}V^{\top}\\ \bar{V}^{\top}\end{bmatrix}=W\Sigma\bar{V}^{\top}$ (1)
with $\Sigma=\operatorname{diag}(\sigma_{q+1},\dots,\sigma_{n})$ and $0<\sigma_{q+1}\leq\dots\leq\sigma_{n}$. Naturally $V,\bar{W}\in\mathbb{R}^{n\times q}$ correspond to the null singular values and $\bar{V},W\in\mathbb{R}^{n\times(n-q)}$ correspond to the non-zero singular values. By the properties of the SVD, the columns of $V$ and $W$ form orthonormal bases for the null and range spaces of $J^{\star}$, while those of $\bar{V}$ and $\bar{W}$ form orthonormal bases their respective orthogonal complement subspaces:
$\ker(J^{\star})$ $=\operatorname{im}(V),$ $\quad\operatorname{range}(J^{\star})$ $=\operatorname{im}(W),$
$\ker(J^{\star})^{\perp}$ $=\operatorname{im}(\bar{V}),$ $\quad\operatorname{range}(J^{\star})^{\perp}$ $=\operatorname{im}(\bar{W}).$ (2)
Consequently, the corresponding orthogonal projections are
@@
$P_{\ker(J^{\star})^{\perp}}$ $=\bar{V}\bar{V}^{\top},$ $\quad P_{\operatorname{range}(J^{\star})^{\perp}}$ $=\bar{W}\bar{W}^{\top}.$ (3)
Using these, the decomposition of the state $x\in\mathbb{R}^{n}$ into $\ker(J^{\star})\oplus\ker(J^{\star})^{\perp}$ can be denoted by
$X\ni x\mapsto(x^{\parallel},x^{\perp}):=(VV^{\top}x,\bar{V}\bar{V}^{\top}x)\in\ker(J^{\star})\times\ker(J^{\star})^{\perp}.$
@@
$x=\Gamma(\alpha,\beta):=\underbrace{V\alpha}_{x^{\parallel}}+\underbrace{\bar{V}\beta}_{x^{\perp}}\quad\text{where}\quad\begin{cases}\alpha=V^{\top}x\in\mathbb{R}^{q}\\ \beta=\bar{V}^{\top}x\in\mathbb{R}^{n-q}\end{cases}$ (4)
where $\alpha$ and $\beta$ represent the reduced coordinates of $x^{\parallel}$ and $x^{\perp}$ corresponding to their components in $\ker(J^{\star})$ and its orthogonal complement $\ker(J^{\star})^{\perp}$, respectively. We employ this coordinate map as a shorthand throughout the analysis.
The Lyapunov–Schmidt reduction involves applying the Implicit Function theorem to obtain a locally valid implicit mapping $(\alpha,p)\mapsto\beta$. In particular, it guarantees the existence of open neighborhoods $\mathcal{N}\big((\alpha_{0},p^{\star})\big)\subset\mathbb{R}^{q}\times\Lambda$ and $\mathcal{N}(\beta_{0})\subset\mathbb{R}^{n-q}$ equipped with implicit map $\varphi:\mathcal{N}\big((\alpha_{0},p^{\star})\big)\to\mathcal{N}(\beta_{0})$ that expresses coordinates along the range of the Jacobian as a function of the coordinates along the null space and the bifurcation parameter, $\beta=\varphi(\alpha,p)$, satisfying
$\varphi(\alpha^{\star},p^{\star})$ $=\beta^{\star},$ (5a)
$\Phi(V\alpha+\bar{V}\varphi(\alpha,p),p)$ $=0\quad\forall\;(\alpha,p)\in\mathcal{N}\big((\alpha^{\star},p^{\star})\big).$ (5b)

### v3

@@ Abstract
Local bifurcation analysis plays a central role in understanding qualitative transitions in networked nonlinear dynamical systems, including dynamic neural network and opinion dynamics models. In this article we establish explicit bounds of validity for the classification of bifurcation diagrams in two classes of continuous-time networked dynamical systems, analogous in structure to the Hopfield and the Firing Rate dynamic neural network models. Our approach leverages recent advances in computing the bounds for the validity of Lyapunov–Schmidt reduction, a reduction method widely employed in nonlinear systems analysis. Using these bounds we rigorously characterize neighbourhoods around bifurcation points where predictions from reduced-order bifurcation equations remain reliable. We further demonstrate how these bounds can be applied to an illustrative family of nonlinear opinion dynamics on $k$-regular graphs, which emerges as a special case of the general framework. These results provide new analytical tools for quantifying the robustness of bifurcation phenomena in dynamics over networked systems and highlight the interplay between network structure and nonlinear dynamical behaviour.
@@ I Introduction
Local bifurcations are points of critical transitions in behaviour of nonlinear dynamical systems that occur when a parameter crosses a threshold, causing the stability of an equilibrium point to change and new features such as new equilibria to emerge. Bifurcations are a common route to multistability in complex systems, and bifurcation analysis is used to characterize the emergence, stability, and basins of attraction of coexisting equilibria as parameters vary.
@@
In this paper we study the region of validity for local bifurcation analysis in two broad classes of nonlinear networked dynamical systems, analogous in their structure to Hopfield and firing rate dynamic neural network models. Systems of this form are commonly encountered as models of dynamic neural networks [14, 3, 23, 2], opinion dynamics and decision-making [7, 4, 12, 6, 1, 17], Lur’e feedback systems [22, 18, 20], gene regulatory networks [11, 21, 8], and other complex systems across application domains. Multistable equilibria are common desirable features in these models, representing meaningful states such as stored associative memories in a neural network or collective decision states in a group of communicating agents. Lyapunov–Schmidt reduction is a common nonlinear analysis step in the local bifurcation analysis of nonlinear networks to classify such multistabilities [10, 17]. A central challenge is that the method is inherently local, yet the size of the neighbourhood in which the reduced bifurcation equations faithfully capture the local bifurcation diagram of the full system is typically unknown. In this work we will bound these regions of validity for the two broad classes of networked dynamical systems.
@@
In our prior work [13], we developed estimates for the domain of validity of reduced-order bifurcation equations for local bifurcations of equilibria derived using Lyapunov–Schmidt reduction. These arise as finite-dimensional equilibrium equations defined on the kernel of the linearized dynamics at the singular point. Our results built directly upon the those established in [16], where explicit bounds of validity were obtained for the application of Implicit Function theorem using tools from functional analysis. In the present work, we build on the general results by specializing the bounds in the context of equilibrium bifurcations in networked dynamical systems. Specifically, we apply this methodology to Hopfield and Firing Rate network models, relating the regions of validity of Lyapunov–Schmidt reduction-based analysis to network structure and choice of smooth nonlinear activation functions in these models.
@@
Our contributions are organized as follows. After introducing notation, relevant lemmas, and key background results on explicit bounds of validity for the Lyapunov–Schmidt reduction in Section II, we present in Section III a simplified general bound on the regions of validity for Lyapunov–Schmidt reduction in an arbitrary nonlinear dynamical system, building on our previous work in [13]. Next, in Section IV we specialize this bound for the Hopfield and Firing Rate network models, presenting simplified computable expressions that leverage the structure of the model equations. Subsequently, in Section V we apply the derived bounds to consensus bifurcations in nonlinear opinion dynamics networks over regular graphs. We show that the resulting bounds involve interpretable quantities related to the structure of the underlying graph. Finally, in Section VI we provide a brief discussion.
@@ II-A Notations and Preliminaries
Let $\mathbb{R}$ denote the set of real numbers and $\mathbb{N}$ the set of positive integers. A mapping $f:U\to V$, where $U\subseteq\mathbb{R}^{m}$ and $V\subseteq\mathbb{R}^{n}$, is said to be of class $C^{\nu}$ for $\nu\in\mathbb{N}$ if it is at least $\nu$-times continuously differentiable. For $f\in C^{\nu}$, we write $D_{y}f(y)$ for the Jacobian matrix of partial derivatives with respect to the variables $y$. For $y_{0}\in\mathbb{R}^{n}$ and $r>0$, the open ball of radius $r$ centered at $y_{0}$ is denoted by $\mathcal{B}(y_{0},r):=\{y\in\mathbb{R}^{n}:\left\lVert{y-y_{0}}\right\rVert<r\}$, where $\left\lVert{\cdot}\right\rVert$ is a vector norm on $\mathbb{R}^{n}$. For $A\in\mathbb{R}^{p\times q}$, we use $\left\lVert{A}\right\rVert$ to denote the spectral matrix norm induced by the same vector norm. For a matrix $M$, we denote by $\sigma_{i}(M)=\lambda_{i}(M^{\top}M)$ its $i$-th largest singular value and by $\sigma_{\min}(M)$ and $\sigma_{\max}(M)$ its smallest and largest singular values respectively. We let $\rho(M)$ denote the spectral radius of $M$. We write $\mathbf{0}_{n},\mathbf{1}_{n}\in\mathbb{R}^{n}$ for the all-zero and all-one column vectors, respectively, and $\mathbf{0}_{p\times q},\mathbf{1}_{p\times q}\in\mathbb{R}^{p\times q}$ for the corresponding matrices.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph with vertex set $\mathcal{V}$ of cardinality $|\mathcal{V}|=n$ and edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. We associate with $\mathcal{G}$ the signed adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $a_{ij}$ is the signed weight of the edge between nodes $i$ and $j$, and $a_{ij}=0$ if no such edge exists. A graph is said to be $k$-regular if every vertex has exactly $k$ neighbours, i.e., the degree of each vertex is equal to $k$. Equivalently, the adjacency matrix $A$ of a $k$-regular graph satisfies $A\mathbf{1}_{n}=k\mathbf{1}_{n}$.
The following lemma provides a convenient matrix norm identity that will be used repeatedly in our analysis.
@@
Let $U\in\mathbb{R}^{n\times k}$ be a matrix with orthonormal columns and let $P=UU^{\top}$ be the orthogonal projection matrix onto $\mathrm{col}(U)$. Then for any $Y\in\mathbb{R}^{n\times m}$, we have $\left\lVert{U^{\top}Y}\right\rVert=\left\lVert{PY}\right\rVert$ and $\left\lVert{YU}\right\rVert=\left\lVert{YP}\right\rVert$.
@@
Since $U^{\top}U=I$, and $P^{2}=P$, from the definition of spectral radius we have $\left\lVert{U^{\top}Y}\right\rVert^{2}=\rho(Y^{\top}UU^{\top}Y)=\rho(Y^{\top}PY)=\rho(Y^{\top}P^{2}Y)=\left\lVert{PY}\right\rVert^{2}.$ Using $P^{\top}=P$, we also have $\left\lVert{YU}\right\rVert=\left\lVert{U^{\top}Y^{\top}}\right\rVert=\left\lVert{PY^{\top}}\right\rVert=\left\lVert{YP}\right\rVert.$
@@ II-B Lyapunov–Schmidt reduction
Lyapunov–Schmidt reduction is a method used to analyse parametrized nonlinear dynamical systems near their singular equilibria, leveraging the Implicit Function theorem. The method decomposes the state space of a dynamical system into the null space of its Jacobian and its orthogonal complement at the singular point. The Implicit Function theorem is then used to eliminate the component orthogonal to the null space in the local bifurcation diagram of the system, effectively projecting the high-dimensional equations describing the equilibria of the model onto the kernel [9, Ch. VII]. The resulting reduced equations capture the structure of the local bifurcation diagram of the original system. We now present a partial summary of the reduction procedure.
Let $X\subset\mathbb{R}^{n}$ and $\Lambda\subset\mathbb{R}^{m}$ be open sets. Consider the parametrized nonlinear dynamical system $\dot{x}=\Phi(x,p)$ where $x\in X$ is the state vector, $p\in\Lambda$ is the parameter vector, and $\Phi:X\times\Lambda\to\mathbb{R}^{n}$ is at least $\mathcal{C}^{2}$, with a bifurcation at some singular point $(x^{\star},p^{\star})\in X\times\Lambda$, i.e. $\Phi(x^{\star},p^{\star})=0$ and $J^{\star}:=D_{x}\Phi(x^{\star},p^{\star})$ is singular (non-invertible) with $q:=\dim\bigl(\ker(J^{\star})\bigr)>0$.
@@
$J^{\star}=\begin{bmatrix}W&\overline{W}\end{bmatrix}\begin{bmatrix}\Sigma&\mathbf{0}_{q\times(n-q)}\\ \mathbf{0}_{(n-q)\times q}&\mathbf{0}_{q\times q}\end{bmatrix}\begin{bmatrix}\overline{V}^{\top}\\ V^{\top}\end{bmatrix}=W\Sigma\overline{V}^{\top}$ (1)
with $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n-q})$ and $\sigma_{1}\geq\dots\geq\sigma_{n-q}>0$. Naturally $V,\overline{W}\in\mathbb{R}^{n\times q}$ correspond to the null singular values and $\overline{V},W\in\mathbb{R}^{n\times(n-q)}$ correspond to the non-zero singular values. By the properties of the SVD, the columns of $V$ and $W$ form orthonormal bases for the null and range spaces of $J^{\star}$, while those of $\overline{V}$ and $\overline{W}$ form orthonormal bases their respective orthogonal complement subspaces. Consequently, the corresponding orthogonal projections are
@@
$P_{\ker(J^{\star})^{\perp}}$ $=\overline{V}\,\overline{V}^{\top},$ $\quad P_{\operatorname{range}(J^{\star})^{\perp}}$ $=\overline{W}\,\overline{W}^{\top}.$ (2)
Using these, the unique orthogonal decomposition of the state $x\in\mathbb{R}^{n}$ into $\ker(J^{\star})\oplus\ker(J^{\star})^{\perp}$ can be denoted by
$X\ni x\mapsto(x^{\parallel},x^{\perp}):=(VV^{\top}x,\overline{V}\,\overline{V}^{\top}x)\in\ker(J^{\star})\times\ker(J^{\star})^{\perp}.$
@@
$x=\Gamma(\alpha,\beta):=\underbrace{V\alpha}_{x^{\parallel}}+\underbrace{\overline{V}\beta}_{x^{\perp}}\quad\text{where}\quad\begin{cases}\alpha=V^{\top}x\in\mathbb{R}^{q}\\ \beta=\overline{V}^{\top}x\in\mathbb{R}^{n-q}\end{cases}$ (3)
where $\alpha$ and $\beta$ are the coordinates of $x^{\parallel}$ and $x^{\perp}$ in the bases $V$ and $V^{\top}$, respectively. We employ this coordinate map as a shorthand throughout the analysis.
The Lyapunov–Schmidt reduction involves applying the Implicit Function theorem to obtain a locally valid implicit mapping $(\alpha,p)\mapsto\beta$. In particular, it guarantees the existence of open neighbourhoods $\mathcal{N}\big((\alpha_{0},p^{\star})\big)\subset\mathbb{R}^{q}\times\Lambda$ and $\mathcal{N}(\beta_{0})\subset\mathbb{R}^{n-q}$ equipped with implicit map $\varphi:\mathcal{N}\big((\alpha_{0},p^{\star})\big)\to\mathcal{N}(\beta_{0})$ that expresses coordinates of the state $x$ along the $\ker(J^{\star})^{\perp}$ as a function of its coordinates along $\ker(J^{\star})$ and the bifurcation parameter, i.e. $\beta=\varphi(\alpha,p)$, satisfying
$\Phi(V\alpha+\overline{V}\varphi(\alpha,p),p)$ $=0\quad\forall\;(\alpha,p)\in\mathcal{N}\big((\alpha^{\star},p^{\star})\big).$ (4)